OpenStudy (anonymous):

Which of the following are solutions to the equation below? Check all that apply. x2 + 10x + 25 = 12 A. x = - - 10 B. x = 2 - 5 C. x = - 10 D. x = -2 - 5 E. x = 12 + F. x = 12 -

OpenStudy (anonymous):

sorry ignore the answer choices

OpenStudy (thomas5267):

Use the quadratic equation. \[ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \] where \(a\) is the coefficient of the term \(x^2\), \(b\) is the coefficient of \(x\) and \(c\) is the constant.

OpenStudy (thomas5267):

Rearrange the equation into this form first. \[ ax^2+bx+c=0 \] Then find a, b and c and use the quadratic equation.

OpenStudy (anonymous):

\[\frac{ -10\pm \sqrt{48} }{ 2 }\]